# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v1

@@ 3.4. Text Classifier Training with Path-Based Data Augmentation
Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a textual entailment model $\mathcal{M}$, a sentence encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
@@
2 for $d_{i}\in\mathcal{D}$ do
3       $\mathbb{C}^{0}_{i}\leftarrow$ use the LLM $\mathcal{G}$ to select core classes from candidates retrieved by $\mathcal{M}$;
@@
5// Corpus-Based Taxonomy Enrichment;
@@
7       $D^{0}_{c}\leftarrow$ a set of roughly classified documents (Eq. 1);
@@
9             $T(c,c_{p})\leftarrow$ top terms ranked by affinity (Eq. 4);
10
11      $T_{c}\leftarrow$ aggregate class-indicative terms with Eq. 5;
12
13// Core Class Refinement with Enriched Taxonomy;
@@ 24 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
25 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$, $\mathcal{D}^{\text{gen}}$, and Eq. 7;
@@ 26 Return $F$;
Table 1. Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR. The best score among zero-shot and weakly-supervised methods is boldfaced. “${}^{\dagger}$” indicates the numbers for these baselines are directly from previous paper (Shen et al., 2021). “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.
Supervision Type Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Zero-Shot ${}^{\dagger}$ 0.4742 0.7144 0.4610 0.6765 0.7871 0.6765
GPT-3.5-turbo 0.5164 0.6807 0.4752 0.4816 0.5328 0.4547
Weakly-Supervised ${}^{\dagger}$ 0.3157 0.5805 0.3115 0.1443 0.2635 0.1443
${}^{\dagger}$ 0.2458 0.5773 0.2517 0.3047 0.5359 0.3048
${}^{\dagger}$ 0.5431 0.7918 0.5414 0.5911 0.7712 0.8621 0.7712 0.8221
${}^{\dagger}$ 0.5934 0.8120 0.5894 0.6332 0.8156 0.8942 0.8156 0.8762
TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
4. Experiments
4.1. Experiment Setup
4.1.1. Datasets
We use two public datasets in different domains for evaluation. (1) Amazon-531 (McAuley and Leskovec, 2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples. (2) DBpedia-298 (Lehmann et al., 2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
4.1.2. Compared Methods
@@
  • Hier-0Shot-TC (Yin et al., 2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
  • GPT-3.5-turbo is a zero-shot approach that queries GPT-3.5-turbo by directly providing all classes in the prompt.
@@
  • Hier-doc2vec (Le and Mikolov, 2014) is a weakly-supervised approach, which first trains document and class representations in the same embedding space, and then iteratively selects the most similar class at each level.
@@
  • WeSHClass (Meng et al., 2019) is a weakly-supervised approach using a set of keywords for each class. It first generates pseudo documents to pretrain text classifiers and then performs self-training.
@@
  • TaxoClass and its variant TaxoClass-NoST (Shen et al., 2021) is a weakly-supervised approach that only uses the class name of each class. It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo-training data. We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as ours.
  • TELEClass is our newly proposed weakly-supervised approach that only uses the class name for each class.
  • Fully-Supervised is a fully-supervised baseline that uses the entire labeled training data to train the text-matching network used in TELEClass.
4.1.3. Evaluation Metrics
Following previous studies (Shen et al., 2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR). See Appendix A for definitions of these metrics.
4.1.4. Implementation Details
We use RoBERTa-large-MNLI as the textual entailment model for core class candidate search. We query GPT-3.5-turbo-0613 for both core class annotation and path-based generation. For taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased. The weight parameter $\alpha$ is set to 0.2, and we select top $k=20$ enriched terms for each class (c.f. Section 3.2). For core class refinement, we get the document and class embeddings using Sentence Transformer all-mpnet-base-v2 (c.f. Section 3.3). We generate $q=5$ documents with path-based generation for each class. The document encoder in the final classifier is initialized with BERT-base-uncased. We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64. We include prompts we used for LLM annotation and generation in Appendix B.
4.2. Experimental Results
Table 1 shows the evaluation results of all the compared methods. We make the following observations. (1) Overall, TELEClass achieves the best performance among the compared strong zero-shot and weakly-supervised baselines. (2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, which is the strongest baseline that also does not use self-training as TELEClass. This further demonstrates the effectiveness of TELEClass. (3) Although LLM (e.g., GPT-3.5-turbo) show superior power in many tasks, naïvely prompting GPT-3.5-turbo in the hierarchical text classification task does not yield promising performance and underperforms strong weakly-supervised text classifier by a large margin. We will conduct a more detailed comparison of LLM prompting for hierarchical text classification in Section 4.4.
Table 2. Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298 datasets. The best score is boldfaced.
Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Gen-Only 0.5151 0.7477 0.5096 0.5357 0.7930 0.9421 0.7930 0.8209
TELEClass-NoEnrich 0.6324 0.8375 0.6263 0.6730 0.6868 0.7726 0.6868 0.7182
TELEClass-NoGen 0.6084 0.8238 0.6024 0.6361 0.8364 0.9146 0.8364 0.8599
TELEClass 0.6330 0.8439 0.6269 0.6664 0.8684 0.9293 0.8684 0.8880
Table 3. Performance comparison of zero-shot LLM prompting. We only report Example-F1 and P@k, because it is not straightforward to get ranking of classes predicted by LLMs for MRR calculation. We also report estimated cost in US dollar (as of the pricing in Feb 2024) and running time in minutes for each method on the entire test set. “${}^{\ddagger}$” indicates that we report the performance on a 1,000-document subset of test data.
Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 Est. Cost Est. Time Example-F1 P@1 P@3 Est. Cost Est. Time
GPT-3.5-turbo 0.5164 0.6807 0.4752 \$60 240 mins 0.4816 0.5328 0.4547 \$80 400 mins
GPT-3.5-turbo (level) 0.6621 0.8574 0.6444 \$20 800 mins 0.6649 0.8301 0.6488 \$60 1,000 mins
${}^{\ddagger}$ 0.6994 0.8220 0.6890 \$800 400 mins 0.6054 0.6520 0.5920 \$2,500 1,000 mins
TELEClass 0.6330 0.8439 0.6269 <\$1 3 mins 0.8684 0.9293 0.8684 <\$1 7 mins
4.3. Ablation Studies
To better understand if each component of TELEClass contributes to the entire method, we additionally conduct ablation studies by removing part of TELEClass and reporting the performance. Table 2 shows the results of the following ablations of TELEClass:
  • Gen-Only only uses the augmented documents produced by path-based LLM generation to train the final classifier.
  • TELEClass-NoEnrich does not perform taxonomy enrichment, so directly uses the initial core classes annotated by LLM plus augmented documents to train the final classifier.
  • TELEClass-NoGen does not use the augmented documents produced by path-based LLM generation, so only uses the refined core classes by taxonomy enrichment and embedding-based matching to train the final classifier.
We find that the full model TELEClass achieves the overall best performance among the compared methods, showing the effectiveness of each of its components. Interestingly, we find that taxonomy enrichment and path-based data generation make different levels of contribution on two data sets: path-based data generation brings more improvement on Amazon-531, while taxonomy enrichment contributes more on DBpedia-298. The reasons could be as follows. Amazon-531 has a large label space and more skewed class distribution. We find that around 11.6% of classes are not selected as core classes in Amazon, while this number for DBpedia is 5.4%. This necessitates data augmentation to provide training data that covers each class in the taxonomy. On the other hand, although DBpedia has a smaller label space, its class names are more subtle to distinguish, which can also be demonstrated by the lower performance of zero-shot LLM prompting on DBpedia compared to Amazon-531 (c.f. GPT-3.5-turbo in Table 1). Therefore, by mining class-indicative terms, TELEClass can enrich each class with more clues to facilitate better classification.
4.4. Comparison with Zero-Shot LLM Prompting
In this section, we further compare TELEClass with zero-shot LLM prompting. Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations. Additionally, we report the estimated cost and time for each method on the entire test set. The cost is reported in US dollar based on the pricing in February 2024. The inference time is reported in minutes, and please be aware that this is just a rough estimation as the actual running time is also dependent on the server condition.
We include the following settings to compare with TELEClass:
  • GPT-3.5-turbo: We include all the classes in the prompt and ask GPT-3.5-turbo model to provide 3 most appropriate classes for a given document.
  • GPT-3.5-turbo (level): We perform level-by-level prompting using GPT-3.5-turbo. Starting from the root node, we ask the model to return one most appropriate class for a given document, and we iteratively prompt the model with the child nodes of the selected node at each level. Notice that this method can only generate a path in the taxonomy, but in the actual multi-label hierarchical classification setting, the true labels may not sit in the same path in the label taxonomy.
  • GPT-4: Similar to the first one, we include all the classes in the prompt and ask GPT-4 to provide 3 most appropriate classes for a given document. Given the limited budget, we only test it on a sampled subset of the test split, which contains 1,000 documents for each dataset.
Table 3 shows the experiment results. We find that TELEClass consistent outperforms all compared methods on DBpedia, while on Amazon, TELEClass underperforms GPT-3.5-turbo (level) and GPT-4 but still being comparable. As for the cost, once trained, TELEClass does not require an additional cost on inference and also has substantially shorter inference time. Prompting LLMs takes longer time and can be prohibitively expensive (e.g., using GPT-4), and the cost will scale up with an increasing size of test data. Also, we find that the level-by-level version of GPT-3.5-turbo consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure. It saves the cost because of the much shorter prompts, but takes longer time due to more queries made per document.
Table 4. Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results. The optimal core class in the true labels is marked with ©.
Dataset Document Core Classes by… True Labels Corr. Enrichment
DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Top Enrichment: national library, central library, collection, volumes…
Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room. It gives her the stability and security she needs. TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Top Enrichment: seat, toilet, shower, safety, handles…
4.5. Case studies
To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass (Shen et al., 2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass. Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table. Overall, we can see that TELEClass’s refined core class is the most accurate. For example, for the first Wikipedia article about a library, TaxoClass selects “village” as the core class, while TELEClass’s initial core class finds a closer one “building” thanks to the power of LLMs. Then, with the enriched classes-indicative features as guidance, TELEClass’s refined core class correctly identifies the optimal core class, which is “library”. In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
5. Related Work
5.1. Weakly-Supervised Text Classification
Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency. Various sources of weak supervision have been explored, including distant supervision (Gabrilovich and Markovitch, 2007; Chang et al., 2008; Song and Roth, 2014) like knowledge bases, keywords (Agichtein and Gravano, 2000; Tao et al., 2018; Meng et al., 2018, 2020; Wang et al., 2021a; Zhang et al., 2021b), and heuristic rules (Ratner et al., 2016; Badene et al., 2019; Shu et al., 2020). Later, extremely weakly-supervised methods are proposed to solely rely on class label names for generating pseudo-labels and training classifiers. LOTClass (Meng et al., 2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords. XClass (Wang et al., 2021a) extracts keywords for creating static class representations through clustering. PIEClass (Zhang et al., 2023a) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training. MEGClass (Kargupta et al., 2023) acquires contextualized sentence representations to capture topical information at the document level. WOTClass (Wang et al., 2023) focuses on mining and ranking class-indicative keywords from generated classes and extracting classes with overlapping keywords.
5.2. Hierarchical Text Classification
A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification. Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches. Local approaches train multiple classifiers for each node or local structures (Banerjee et al., 2019; Liu et al., 2005; Wehrmann et al., 2018). Global approaches, learn hierarchy structure into a single classifier through recursive regularization (Gopal and Yang, 2013), a graph neural network (GNN)-based encoder (Peng et al., 2018; Huang et al., 2019; Zhou et al., 2020; Shen et al., 2021), or a joint document label embedding space (Chen et al., 2021).
Weak supervision is also studied for hierarchical text classification. WeSHClass (Meng et al., 2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training. TaxoClass (Shen et al., 2021) follows the same setting as ours which uses the sole class name of each class as the only supervision. It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier. Additionally, MATCH (Zhang et al., 2021c) and HiMeCat (Zhang et al., 2021a) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
5.3. LLMs as Generators and Annotators
Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators (Ding et al., 2023). For data generation, few-shot examples (Wang et al., 2021b) or class-conditioned prompts (Meng et al., 2022) are explored for LLM generation and the generated data can be used as pseudo-training data to further fine-tune a small model as the final classifier (Ye et al., 2022). Recently, Yu et al. (2023) propose an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation. For data annotation, previous works utilize LLMs for unsupervised annotation (Feng et al., 2021), Chain-of-Thought annotation with explanation generation (He et al., 2023), and active annotation (Zhang et al., 2023b).
6. Conclusion and Future Work
In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions. First, we enrich the input label taxonomy with more class-indicative topical terms for each class, which can serve as additional features to understand the classes and facilitate classification. Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation. On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies. We also conduct a comparative analysis on performance and cost for zero-shot LLM prompting for the hierarchical text classification task. For future works, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing. We also plan to explore how to extend TELEClass into a more complicated label space, such as an extremely large space with millions of labels.
Appendix A Evaluation Metrics
@@
  • Example-F1 (Sørensen, 1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
@@
    (10) $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$
Appendix B Prompts for LLM
@@
    Instruction: You will be provided with an Amazon product review, and your task is to select its product types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]
@@
    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and your task is to select its types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]

### v2

@@ Appendix A Pseudo code of TELEClass
Input: A corpus $\mathcal{D}$, a label taxonomy $\mathcal{T}$, a pretrained text encoder $\mathcal{S}$, an LLM $\mathcal{G}$.
@@
2 for $c\in\mathcal{C}$ do
3       $T^{\text{LLM}}_{c}\leftarrow$ use $\mathcal{G}$ to enrich $c$ with key terms;
@@
5for $d_{i}\in\mathcal{D}$ do
6       $\mathbb{C}^{0}_{i}\leftarrow$ use $\mathcal{G}$ to select core classes from candidates retrieved using $\mathcal{S}$ and $T^{\text{LLM}}_{c}$;
7
8// Corpus-Based Taxonomy Enrichment;
@@
10       $D^{0}_{c}\leftarrow$ a set of roughly classified documents;
@@
12             $T(c,c_{p})\leftarrow$ top terms ranked by affinity;
13
14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms;
15
16// Core Class Refinement with Enriched Taxonomy;
@@ 27 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$;
@@ 29 Return $F$;
Appendix B Experiment Setup
B.1 Compared Methods
@@
  • Hier-0Shot-TC Yin et al. (2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
@@
  • Hier-doc2vec Le and Mikolov (2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
@@
  • WeSHClass Meng et al. (2019) is a weakly-supervised approach using a set of keywords for each class. It first generates pseudo documents to pretrain text classifiers and then performs self-training.
@@
  • TaxoClass Shen et al. (2021) is a weakly-supervised approach that only uses the class name of each class. It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data. We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
B.2 Evaluation Metrics
@@
  • Example-F1 Sørensen (1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
@@
    $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$ (5)
B.3 Implementation Details
We use Sentence Transformer Reimers and Gurevych (2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3. We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation. For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased Devlin et al. (2019), we select top $k=20$ enriched terms for each class (c.f. Section 3.2). We generate $q=5$ documents with path-based generation for each class. The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines. We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64. The experiments are run on one NVIDIA RTX A6000 GPU.
Appendix C Prompts for LLM
  • LLM-based enrichment for Amazon-531
    Instruction: [Target Class] is a product class in Amazon and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
  • LLM-based enrichment for DBPedia-298
    Instruction: [Target Class] is a article category of Wikipedia articles and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
@@ Appendix C Prompts for LLM
    Instruction: You will be provided with an Amazon product review, and please select its product types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]
@@
    Instruction: You will be provided with a Wikipedia article describing an entity at the beginning, and please select its types from the following categories: [Candidate Classes]. Just give the category names as shown in the provided list. Query: [Document]